OpenStudy (unicornpoopcookies):

solve using the principle of zero products (x+3)(5x-4)=0

Nnesha (nnesha):

a*b=0 a=0 or b=0 set both parentheses equal to zero and then solve for x

OpenStudy (unicornpoopcookies):

I wanted the answer but thanks.

OpenStudy (kkutie7):

\[x+3=0\rightarrow x=-3\] \[5x-4=0\rightarrow 5x=4\rightarrow x=\frac{4}{5}\]